OpenStudy (anonymous):

can someone help me find the solution of •g(x) = x2 +6x + 1

OpenStudy (jhannybean):

g(x) = x\(^2\)+6x+1 I believe this method involves completing the square. Are you familiar with that?

OpenStudy (anonymous):

yeah, I know it's irrational though because it has 2 real roots

OpenStudy (jhannybean):

Mmhmm, so let's begin by the completing the square method.

OpenStudy (anonymous):

can you walk me through it please :3 I'm not exactly familiar with that that's why I asked :l it's my last question of the problem though

OpenStudy (jhannybean):

What you want to do first, is set your equation = 0. \[g(x) = x^2 +6x+1=0\]

OpenStudy (jhannybean):

Now subtract both sides of the equation by -1 to isolate your x values. \[x^2+6x=-1\]

OpenStudy (anonymous):

so -1 =x^2+6x?

OpenStudy (jhannybean):

Now we need to complete the square. You can do this by taking your value of b, dividing by 2, and squaring it. This will produce your `c` value, which puts the left hand side of your equation into quadratic form, `ax+by+c`

OpenStudy (jhannybean):

Yes, you're on track :)

OpenStudy (anonymous):

so c would = 9?

OpenStudy (jhannybean):

So now we will have \[c=\left(\frac{b}{2}\right)^2 = \left(\frac{6}{2}\right)^2= (3)^2 =9\]\[x^2+6x\color{blue}{+9} = -1\color{blue}{+9}\]

OpenStudy (jhannybean):

Yep, C = 9.

OpenStudy (anonymous):

sweet thank you so much :)

OpenStudy (jhannybean):

Now we see that the left hand side of the equation is a perfect square, therfore we can reduce it to \((x+c)^2 \implies (x+3)^2\)\[(x+3)^2 = 8\]

OpenStudy (jhannybean):

You're following, yes? :)

OpenStudy (anonymous):

yes :)

OpenStudy (jhannybean):

Awesome. To solve for x, we take the square root of both sides. \[\sqrt{(x+3)^2} = \pm \sqrt{8}\]

OpenStudy (jhannybean):

This will help us undo the square that is bounding the x, and help us solve for it. \[x+3=\pm \sqrt{8}\]

OpenStudy (jhannybean):

Now just subtract `-3` from both sides of your equation and you will have solved for x :D \[x= -3 \pm \sqrt{8}\]

OpenStudy (anonymous):

take the square root of 8, then subtract for 3?

OpenStudy (jhannybean):

Precisely.

OpenStudy (anonymous):

ouch spoke to fast, so you don't take the square root of 8?

OpenStudy (jhannybean):

depends how simplified you want it to be. \[\sqrt{8} \iff 2\sqrt{2}\]

OpenStudy (jhannybean):

So you would then have \[x=-3\pm 2\sqrt{2}\]

OpenStudy (anonymous):

this makes a lot of sense, thanks :)

OpenStudy (jhannybean):

Awesome :) No problem.